Solve for x in log form
4^2X-4^X-42=0
answer: log(base 4)7 = X

Answer by stanbon(75887)

Solve for x in log form
4^2X-4^X-42=0
---------
Let 4^x = w
---------
Substitute to get:
w^2 - w - 42 = 0
(w-7)(w+6) = 0
w = 7 or w = -6
---------------------
Substitute to solve for "x":
4^x = 7 or 4^x = -6
No power of 4 can equal -6 so the only solution is that of 4^x=7
Takethe log of both sides to get:
xlog(4) = log(7)
x = log(7)/log(4)
x = 1.4036775...
=========================
Cheers,
Stan H.